OpenStudy (anonymous):

Find all the integers b for which a^2+ab-50 can be factored.

OpenStudy (mayankdevnani):

Need two numbers whose product is -50 and whose sum is "b": Products are 1*-50 with sum = b = -49 -1*50 with sum = b = 49 --- 2*-25 with sum = b = -23 -2*25 with sum = b = 23 ---- -5*10 with sum = b = 5 5*-10 with sum = b = -5 Those are the possible values of b if you limit b to being an integer.

OpenStudy (mayankdevnani):

@stupidinmath

OpenStudy (anonymous):

ohh. that makes sense// thanks

OpenStudy (mayankdevnani):

welcome